# On a Continuum Model for Random Genetic Drift: A Dynamical Boundary Condition Approach

### v2

@@ Abstract.
We propose a new continuum model for a random genetic drift problem by employing a dynamical boundary condition approach. The new model, which can be viewed as a regularized Kimura equation, admits a continuous solution. The existence and uniqueness of the strong solution of the regularized system are shown. We present some numerical results for the regularized model, which indicate that the model can capture the main features of random genetic drift, such as gene fixation and conservation of the first moment.
@@ 1. Introduction
Here, the variable $x$ is the fraction of the focal allele $A_{1}$ in the population and $(1-x)$ is that of $A_{2}$, the function $\rho(x,t)$ is the probability of finding a relative composition $x\in(0,1)$ of gene $A_{1}$ at time $t$. Although (1) is a linear PDE of $\rho(x,t)$, due to the degeneracy of $x(1-x)$ at the boundary $x=0$ and $x=1$ [12, 13], even the appropriate boundary conditions and solution space of the Kimura equation are unclear [16]. In the 1950s, Kolmogorov suggested that equation (1) is only reasonable for $x$ not too close to 00 and $1$ [23].
@@
The measure-valued solution (6) imposed a difficulty to study Kimura equation both numerically [11] and theoretically [3]. For instance, although a lot of different methods for the Kimura equation have been developed, ranging from finite volume method [33, 34], finite Lagrangian methods [11], an optimal mass transportation method [2], and SDE–based simulation methods [7, 20], it is difficult to get a good numerical approximation to (6), which includes the Dirac delta function, and accurately capture the dynamics of $a(t)$ and $b(t)$.
The purpose of this paper is to propose a new continuum model for random genetic drift by incorporating a dynamical boundary approach. The key idea is to only consider the Kimura equation on $(\delta,1-\delta)$ for a given small $\delta>0$, and use a dynamical boundary condition to describe the jump process between bulk and surface, i.e., the flux between $\rho(x,t)$ and $a(t)$/$b(t)$. The new model is given by
(8) $\begin{cases}&\rho_{t}=\partial_{xx}(x(1-x)\rho),\quad x\in(\delta,1-\delta),\\ &\partial_{x}(x(1-x)\rho)\Big{|}_{x=\delta}=a^{\prime}(t),\quad\partial_{x}(x(1-x)\rho)\Big{|}_{x=1-\delta}=-b^{\prime}(t)\\ &a^{\prime}(t)=-((\epsilon a)-\rho(\delta,t))\\ &b^{\prime}(t)=-((\epsilon b)-\rho(1-\delta,t)).\\ \end{cases}$
@@
The rest of this paper is organized as follows. In Section 2, we introduce some background, including the Wright-Fisher model, the variational structure of Kimura equation, and the dynamical boundary approach for generalized diffusions. The new continuum model is presented in Section 3. The existence and uniqueness of the strong solution of the regularized system is shown in Section 4. Finally, we perform a numerical study on the regularized system, and demonstrate the effects of $\delta$ and $\epsilon$, as well as the ability of the new model in capturing the key feature of the original Kimura equation.
@@ 2.1. From a Wright-Fisher model to the Kimura equation
which is known as the pure drift case. Notice that in the pure drift case, the jump rate becomes 0 if $X_{k}=0$ or $X_{k}=1$, which are two absorbing states in the system. There is no possibility of fixations if $u>0$ and $v>0$.
@@ 2.2. Formal Variational Structure
Combining the velocity equation with the kinematics (16), we can recover the original Kimura equation. The variational structure (19) naturally leads to the Lagrangian algorithms developed in [11, 2].
@@
The distance (22) is known as Wasserstein-Shahshahani distance [4, 3].
@@
2.3. Dynamical boundary condition
As suggested in [3], a natural attempt at compensating for singularities is to relax the boundary condition by taking into account of the bulk/surface interaction.
Before we present the application of the dynamical boundary condition approach to the Kimura model, we first briefly review the dynamical boundary approach for generalized diffusions in this subsection. Consider a bounded domain $\Omega$, and let $\partial\Omega$ be the boundary of $\Omega$. Classical PDE models on $\Omega$ often impose Dirichlet, Neumann, and Robin boundary conditions for the physical variable $\rho$. The basic fundamental assumption behind this is that $\rho$ is regular enough (e.g. $\rho\in C^{m}(\bar{\Omega})$ for some $m$), allowing $\rho|_{\partial\Omega}$ to be defined as the trace of $\rho$, and the boundary condition to be specified in terms of $\rho|_{\Gamma}$. However, as in the classical Kimura equation, it may not be always possible to take the trace.
@@ 2.3. Dynamical boundary condition
$\rho\mathbf{u}\cdot{\bm{\nu}}=R_{t},\quad\sigma_{t}+\nabla_{\partial\Omega}\cdot(\sigma{\bm{v}})=R_{t},~{}~{}\bm{x}\in\partial\Omega$
@@
In general, systems with dynamical boundary condition can be modeled through an energy-dissipation law
(25) $\frac{\mathrm{d}}{\mathrm{d}t}\left(\mathcal{F}_{b}(\rho)+\mathcal{F}_{s}(\sigma)\right)=-\left(\int_{\Omega}\eta_{b}(\rho)|\mathbf{u}|^{2}\mathrm{d}\bm{x}+\int_{\partial\Omega}\eta_{s}(\sigma)|{\bm{v}}|^{2}+\dot{R}\Gamma(R,\dot{R})\mathrm{d}S\right)$
Here $\mathcal{F}_{b}(\rho)$ and $\mathcal{F}_{s}(\sigma)$ are free energies in the bulk and surface respectively, $\eta_{b}(\rho)>0$ and $\eta_{s}(\sigma)>0$ are friction coefficients for bulk and surface diffusions, $\dot{R}\Gamma(R,\dot{R})\geq 0$ is the dissipation due to the bulk/surface interaction, which in general is non-quadratic in terms of $\dot{R}$ [30]. The variational procedures lead to the force balance equations for the mechanical and chemical parts (see [30]) for details)
(26) $\begin{cases}&\eta_{b}(\rho)\mathbf{u}=-\rho\nabla\mu_{b}(\rho),\quad\mu_{b}=\frac{\delta\mathcal{F}_{b}}{\delta\rho}\quad\bm{x}\in\Omega\\ &\eta_{s}(\sigma){\bm{v}}=-\sigma\nabla_{\Gamma}\mu_{s}(\sigma),\quad\mu_{s}=\frac{\delta\mathcal{F}_{s}}{\delta\sigma},\quad\bm{x}\in\Gamma\\ &\Gamma(R,\dot{R})=-(\mu_{s}(\sigma)-\mu_{b}(\rho)),\quad\bm{x}\in\Gamma\\ \end{cases}$
@@ 3. A regularized Kimura equation
In this section, we propose a regularized Kimura equation by applying a dynamical boundary condition approach in this specific one dimensional setting.
@@ 3.1. Model derivation
The key idea is to only consider the Kimura equation on $(\delta,1-\delta)$ for a given small $\delta>0$, and use a dynamical boundary condition to describe the fixation on the boundary. Specially, for a given $\delta>0$, let $\rho(x,t)$ represent the probability that the gene frequency is equal to $x\in(\delta,1-\delta)$. We denote the probabilities of the gene frequency being $x=0$ and $x=1$ as $a(t)$ and $b(t)$, respectively. Due to the conservation of mass, we have
@@
where $G_{0}(a)$ and $G_{1}(b)$ are free energy on the boundary, and $\Gamma_{0}(R_{0},\dot{R}_{0})$ and $\Gamma_{1}(R_{1},\dot{R}_{1})$ are dissipations on due to the jump between bulk and surface. The energy-dissipation law in the bulk region ($x\in(\delta,1-\delta)$) is exactly the same as that in the original Kimura equation (15). The remaining question is how to choose $G_{i}(i=0,1)$ and $\Gamma_{i}(i=0,1)$ in order to capture the qualitative behavior of the original Kimura equation. In the current work we take
@@
Here, $\eta(\epsilon)>0$ determines the boundary-surface equilibrium , $\kappa(\epsilon)$ is the reaction rate from surface to bulk. In the current study, we take $\kappa(\epsilon)=\epsilon$ and $\eta(\epsilon)=\frac{\epsilon}{\gamma}$ with $\gamma>0$ being a constant. We’ll explore other choices in future work.
@@
$\ln\left(\frac{\dot{R}_{1}}{\epsilon b}+1\right)=-(\ln(\tfrac{\epsilon}{\gamma}b)-\ln\rho(1-\delta,t))$
@@
(35) $\begin{cases}&\partial_{t}\rho=-\partial_{x}(\rho u),\quad x\in[\delta,1-\delta]\\ &\rho u=-\partial_{x}(x(1-x)\rho),\quad x\in(\delta,1-\delta),\\ &\rho u(\delta,t)=-a^{\prime}(t),\quad\rho u(1-\delta,t)=b^{\prime}(t)\\ &a^{\prime}(t)=\gamma\rho(\delta,t)-\epsilon a\\ &b^{\prime}(t)=\gamma\rho(1-\delta,t)-\epsilon b\\ \end{cases}$
@@
with the initial condition $a(0)=0$. As a consequence
(38) $a^{\prime}(t)=-\epsilon e^{-\epsilon t}\int_{0}^{t}\gamma e^{\epsilon s}\rho(\delta,s)\mathrm{d}s+\gamma\rho(\delta,t)$
@@
(40) $\begin{cases}&\partial_{t}\rho=\partial_{xx}(x(1-x)\rho),\quad x\in(\delta,1-\delta)\\ &\partial_{x}(x(1-x)\rho)|_{x=\delta}=\gamma\rho(\delta,t),\quad\partial_{x}(x(1-x)\rho)|_{x=1-\delta}=-\gamma\rho(1-\delta,t)\\ \end{cases}$
which can be viewed as a closed equation on $(\delta,1-\delta)$ with Robin boundary condition. The system can be interpreted as weighted $L^{2}$-type gradient flow
(41) $\frac{\mathrm{d}}{\mathrm{d}t}\left(\int_{\delta}^{1-\delta}|\partial_{x}(x(1-x)\rho|^{2}\mathrm{d}x+\gamma\delta(1-\delta)(|\rho(\delta,t)|^{2}+|\rho(1-\delta,t)|^{2})\right)=-\int_{\delta}^{1-\delta}x(1-x)|\rho_{t}|^{2}\mathrm{d}x$
@@
(44) $\tilde{\rho}(x,t)=\begin{cases}&a(t)/\delta,\quad x\in[0,\delta)\\ &\rho(x.t),\quad x\in[\delta,1-\delta]\\ &b(t)/\delta,\quad x\in(1-\delta,1]\\ \end{cases}$
@@
$=-(x(1-x)\rho)\Big{|}_{\delta}^{1-\delta}+\frac{\delta}{2}(b^{\prime}(t)-a^{\prime}(t))=-\delta(1-\delta-\tfrac{\gamma}{2})(\rho(1-\delta)-\rho(\delta))+\frac{\epsilon\delta}{2}(a(t)-b(t))$
@@ 3.2. Equilibrium of the regularized model
Hence, for fixed $\delta$, by letting $\epsilon\rightarrow 0$, the equilibrium solution $\rho^{\infty}$ goes to 00.
@@ 4. Existence of Solutions: Operator Approach
Assume that the parameters $\delta,\,\epsilon>0$ and let the initial data satisfy $\rho_{0}\in C(\delta,1-\delta)$ and let $|a_{0}|,\,|b_{0}|<\infty$ be bounded. Then, for all times $T>0$ there exists a unique classical solution $\rho$ to equations (54)-(56) such that
@@
which is a contradiction. Hence, $W\geq 0$ in $[\delta,1-\delta]\times[0,T]$, which implies that $w\geq 0$.
@@ 5.1. Numerical schemes
(59) $a(t)+b(t)+h(\frac{1}{2}\rho_{0}(t)+\sum_{i=1}^{N-1}\rho_{i}(t)+\frac{1}{2}\rho_{N}(t))=1,\quad\forall t$
@@
(62) $\begin{cases}&\frac{\mathrm{d}a}{\mathrm{d}t}=(\rho_{0}-\epsilon a)\\ &\frac{\mathrm{d}\rho_{0}}{\mathrm{d}t}=\frac{2}{h}(-(\rho u)_{1/2}+(\epsilon a-\rho_{0})\\ &\frac{\mathrm{d}\rho_{i}}{\mathrm{d}t}=\frac{1}{h}(-(\rho u)_{i+1/2}+(\rho u)_{i-1/2}),\quad i=1,\ldots N-1\\ &\frac{\mathrm{d}\rho_{N}}{\mathrm{d}t}=\frac{2}{h}((\epsilon b-\rho_{N})+(\rho u)_{N-1/2})\\ &\frac{\mathrm{d}b}{\mathrm{d}t}=(\rho_{N}-\epsilon b)\ ,\\ \end{cases}$
@@ 5.2. Numerical results
Fig. 3(a) shows the numerical solution $\rho(x,t)$ for $\delta=0.001$ and $\epsilon=0.001$ at $t=10$, with $h=10^{-4}$ and $\tau=10^{-4}$. The time evolution of $a(t)$ and $b(t)$ is shown in Fig. 3 (b). The evolution of discrete energy and the first moment are shown in Fig. 3 (c) and (d). In this case, we can also observe the energy stability. Although the absolute difference of the first moment $M(t)$ from its initial value $M_{0}$ is not as small as in the uniform case, it is still small enough to indicate that the first moment is much smaller than $O(\delta)$, which is consistent with the theoretical analysis.

### v3

# On a Continuum Model for Random Genetic Drift: A Dynamic Boundary Condition Approach

@@ Abstract.
We propose a new continuum model for random genetic drift by employing a dynamic boundary condition approach. The model can be viewed as a regularized version of the Kimura equation and admits a continuous solution. We establish the existence and uniqueness of a strong solution to the regularized system. Numerical experiments illustrate that, for sufficiently small regularization parameters, the model can capture key phenomena of the original Kimura equation, such as gene fixation and conservation of the first moment.
@@ 1. Introduction
Here, the variable $x$ is the fraction of the focal allele $A_{1}$ in the population and (1x)(1-x) is that of $A_{2}$, the function $\rho(x,t)$ is the probability of finding a relative composition $x\in(0,1)$ of gene $A_{1}$ at time $t$. Although (1) is a linear PDE of $\rho(x,t)$, due to the degeneracy of x(1x)x(1-x) at the boundary $x=0$ and $x=1$ [12, 13], even the appropriate boundary conditions and solution space of the Kimura equation are unclear [16]. In the 1950s, Kolmogorov suggested that equation (1) is only reasonable for $x$ not too close to 0 and $1$ [23].
@@
The measure-valued solution (6) impose a difficulty to study the Kimura equation both numerically [11] and theoretically [3]. For instance, although a lot of numerical methods for the Kimura equation have been developed, ranging from finite volume method [34, 35], finite Lagrangian methods [11], an optimal mass transportation method [2], and SDE–based simulation methods [7, 20], it is difficult to get a good numerical approximation to (6), which includes the Dirac delta function, and accurately capture the dynamics of $a(t)$ and $b(t)$.
The purpose of this paper is to propose a new continuum model for random genetic drift by incorporating a dynamic boundary approach. The key idea is to only consider the Kimura equation on $(\delta,1-\delta)$ for a given small $\delta>0$, and use a dynamic boundary condition to describe the jump process between bulk and surface, i.e., the flux between $\rho(x,t)$ and the the boundary states $a(t)$ and $b(t)$. The new model is given by
[…]
@@
The rest of this paper is organized as follows. In Section 2, we introduce some background, including the Wright-Fisher model, the variational structure of Kimura equation, and the dynamic boundary approach for generalized diffusions. The new continuum model is presented in Section 3. The existence and uniqueness of the strong solution of the regularized system is shown in Section 4. Finally, we perform a numerical study on the regularized system, and demonstrate the effects of $\delta$ and $\epsilon$, as well as the ability of the new model in capturing the key feature of the original Kimura equation.
@@ 2.1. From a Wright-Fisher model to the Kimura equation
which is known as the pure drift case. Notice that in the pure drift case, the jump rate becomes 0 if $X_{k}=0$ or $X_{k}=1$, which are two absorbing states in the system. These states represent allele fixation, where one allele becomes fixed (frequency 1) and the other lost (frequency 0). The system does not permit fixation if $u>0$ and $v>0$.
@@ 2.2. Formal Variational Structure
Combining the velocity equation with the kinematics (16), we can recover the original Kimura equation. The variational structure (19) naturally leads to the Lagrangian algorithms for the Kimura equation developed in [11, 2].
@@
The distance (22) is known as Wasserstein-Shahshahani distance [2, 4, 3].
@@
2.3. Dynamic boundary condition
As suggested in [3], a natural attempt at compensating for singularities is to relax the boundary condition by taking into account of the bulk/surface interaction. This leads to the use of dynamic boundary conditions [22, 26], in which the boundary values evolve according to their own differential equations, often coupled with the interior dynamics.
Before we present the application of the dynamic boundary condition approach to the Kimura model, we first briefly review the dynamic boundary approach for generalized diffusions in this subsection. Consider a bounded domain $\Omega$, and let $\partial\Omega$ be the boundary of $\Omega$. Classical PDE models on $\Omega$ often impose Dirichlet, Neumann, and Robin boundary conditions for the physical variable $\rho$. The basic fundamental assumption behind this is that $\rho$ is regular enough (e.g. $\rho\in C^{m}(\bar{\Omega})$ for some $m$), allowing $\rho|_{\partial\Omega}$ to be defined as the trace of $\rho$, and the boundary condition to be specified in terms of $\rho|_{\Gamma}$. However, as in the classical Kimura equation, it may not be always possible to take the trace.
@@ 2.3. Dynamic boundary condition
$\rho\mathbf{u}\cdot{\bm{\nu}}=\dot{R},\quad\sigma_{t}+\nabla_{\partial\Omega}\cdot(\sigma{\bm{v}})=\dot{R},~~\bm{x}\in\partial\Omega$
@@
In general, systems with dynamic boundary condition can be modeled through an energy-dissipation law
(25) $\frac{\mathrm{d}}{\mathrm{d}t}\left(\mathcal{F}_{b}(\rho)+\mathcal{F}_{s}(\sigma)\right)=-\left(\int_{\Omega}\eta_{b}(\rho)|\mathbf{u}|^{2}\mathrm{d}\bm{x}+\int_{\partial\Omega}\eta_{s}(\sigma)|{\bm{v}}|^{2}+\dot{R}\Gamma(R,\dot{R})\mathrm{d}S\right)\ .$
Here, $\mathcal{F}_{b}(\rho)$ and $\mathcal{F}_{s}(\sigma)$ are free energies in the bulk and surface respectively, $\eta_{b}(\rho)>0$ and $\eta_{s}(\sigma)>0$ are friction coefficients for bulk and surface diffusions, $\dot{R}\Gamma(R,\dot{R})\geq 0$ is the dissipation due to the bulk/surface interaction, which in general is non-quadratic in terms of $\dot{R}$ [31]. The variational procedures lead to the force balance equations for the mechanical and chemical parts (see [31] for details)
[…]
@@ 3. A regularized Kimura equation
In this section, we propose a regularized Kimura equation by applying a dynamic boundary condition approach in this specific one dimensional setting.
@@ 3.1. Model derivation
The key idea is to only consider the Kimura equation on $(\delta,1-\delta)$ for a given small $\delta>0$, and use a dynamic boundary condition to describe the fixation on the boundary. Specially, for a given $\delta>0$, let $\rho(x,t)$ represent the probability that the gene frequency is equal to $x\in(\delta,1-\delta)$. We denote the probabilities of the gene frequency being $x=0$ and $x=1$ as $a(t)$ and $b(t)$, respectively. Due to the conservation of mass, we have
@@
where $G_{0}(a)$ and $G_{1}(b)$ are free energy on the boundary, and $\Gamma_{0}(R_{0},\dot{R}_{0})$ and $\Gamma_{1}(R_{1},\dot{R}_{1})$ are dissipations on due to the jump between bulk and surface. The energy-dissipation law in the bulk region ($x\in(\delta,1-\delta)$) is exactly the same as that in the original Kimura equation (15). The remaining question is how to choose $G_{i}(i=0,1)$ and $\Gamma_{i}(i=0,1)$ in order to capture the qualitative behavior of the original Kimura equation. In the current work, we take
@@
Here, $\eta(\epsilon)>0$ determines the boundary-surface equilibrium , $\kappa(\epsilon)$ is the reaction rate from surface to bulk. Moreover, we take $\kappa(\epsilon)=\epsilon$ and $\eta(\epsilon)=\frac{\epsilon}{\gamma}$ with $\gamma>0$ being a constant. We’ll explore other choices in future work.
@@
$\ln\left(\frac{\dot{R}_{1}}{\epsilon b}+1\right)=-(\ln(\tfrac{\epsilon}{\gamma}b)-\ln\rho(1-\delta,t))\ .$
@@
[…]
@@
if the initial condition $a(0)=0$. As a consequence,
(38) $a^{\prime}(t)=-\epsilon e^{-\epsilon t}\int_{0}^{t}\gamma e^{\epsilon s}\rho(\delta,s)\mathrm{d}s+\gamma\rho(\delta,t)\ .$
@@
[…]
which can be viewed as a closed equation on $(\delta,1-\delta)$ with Robin boundary condition. Although the energy-dissipation law (29) with (30) is no longer valid, the system can be interpreted as weighted $L^{2}$-type gradient flow
(41) $\frac{\mathrm{d}}{\mathrm{d}t}\left(\int_{\delta}^{1-\delta}|\partial_{x}(x(1-x)\rho)|^{2}\mathrm{d}x+\gamma\delta(1-\delta)(|\rho(\delta,t)|^{2}+|\rho(1-\delta,t)|^{2})\right)=-\int_{\delta}^{1-\delta}x(1-x)|\rho_{t}|^{2}\mathrm{d}x$
@@
[…]
@@
$=-(x(1-x)\rho)\Big{|}_{\delta}^{1-\delta}+\frac{\delta}{2}(b^{\prime}(t)-a^{\prime}(t))=-\delta(1-\delta-\tfrac{\gamma}{2})(\rho(1-\delta)-\rho(\delta))+\frac{\epsilon\delta}{2}(a(t)-b(t))\ .$
@@ 3.2. Equilibrium of the regularized model
Hence, for fixed $\delta$, by letting $\epsilon\rightarrow 0$, the equilibrium solution $\rho^{\infty}$ goes to 0.
@@ 4. Existence of Solutions: Operator Approach
Assume that the parameters $\delta,\,\epsilon>0$ and let the initial data satisfy $\rho_{0}\in C(\delta,1-\delta)$ and let $|a_{0}|,\,|b_{0}|<\infty$ be bounded. Then, for all times $T>0$, there exists a unique classical solution $\rho$ to equations (54)-(56) such that
@@
which is a contradiction. Hence, $W\geq 0$ in $\delta,1-\delta]\times[0,T]$, which implies that $w\geq 0$.
@@ 5.1. Numerical schemes
(59) $a(t)+b(t)+h\left(\frac{1}{2}\rho_{0}(t)+\sum_{i=1}^{N-1}\rho_{i}(t)+\frac{1}{2}\rho_{N}(t)\right)=1,\quad\forall t$
@@
[…]
@@ 5.2. Numerical results
Fig. 3(a) shows the numerical solution $\rho(x,t)$ for $\delta=0.001$ and $\epsilon=0.001$ at $t=10$, with $h=10^{-4}$ and $\tau=10^{-4}$. The time evolution of $a(t)$ and $b(t)$ is shown in Fig. 3 (b). The evolution of discrete energy and the first moment are shown in Fig. 3 (c) and (d). In this case, we can also observe the energy stability, and the first moment is almost a constant.